9. The red super-giant star, Betelgeuse, has a radius of8.21xI m. The spherical surface behaves as a blackbody radiator. If the surface temperature is 3500 K, what is the rate at which energy is radiated from the star? a) 721 1031 w(b210 w (c) 7.7 10 w (d) 2.0 10 w (e) 1.9-100 w 10. The coefficient of linear expansion of a certain solid is 71 C. Assuming this solid behaves like most solids, what is its coefficient of volume expansion? (a) 2.65x10%C (b) 21x 106 C (c) 21x10C (d) 343×104, C (e) 343x1036/C 11. Suppose a railroad track is made of individual rails of steel that are 24 m in length. By what length would these rails change between a cold day when the temperature is-10°C & a hot day at 40°C (a) 1.15 cm (b) 1.44 cm (c) 8.64 em (d) 2.88 cm (e) 0.864 cm
Answer #1

9. given radius r = 8.21*10^11 m

surface temperature, T = 3500 K

rate of energy radiation from the star = r

now from stefan's law

r = sigma*A*e*T^4

sigma = stefans constat = 5.67*10^-8

A = 4pi*r^2

T = 3500 K

e = 1 ( for a blackbody)

hence

r = 72.069484*10^30 W

hence option a) 7.21*10^31 W